OpenStudy (anonymous):

factor the quadratic equation 0=a^2-4a+3

OpenStudy (anonymous):

break the middle term down into a^3 - a - 3a + 3 = 0 so it's easier to factor out

OpenStudy (anonymous):

\[a^2 - 4a + 3 = 0\] \[a^2 - a - a + 3 = 0\] Take out common factors and then solve..

OpenStudy (callisto):

a^2-4a+3 = 0 a^2 - a - 3a + 3 = 0 Factor the left side first.

OpenStudy (anonymous):

(x-1)(x-3) once you factorise it

OpenStudy (anonymous):

soryy its (a-1)(a-3)

OpenStudy (anonymous):

Thanks for all the help I really appreciate it